# A note on subgroups whose quotients are quasi-lines

### v1

@@ 2 Cube complexes and facing triples
First note that $C$ is proper by [Hagen, Corollary 3.9]. In particular, by [Caprace-Sageev, Lemma 4.3], the isometry $g$ is parallel to $\operatorname{\mathfrak{h}}$ if and only if there is a power of $g$ which stabilises $\operatorname{\mathfrak{h}}$. Moreover, it follows from Lemma 2.2 that, if $g$ is peripheral to $\operatorname{\mathfrak{h}}$, then there is some $n\in\mathbb{N}$ such that $\{\operatorname{\mathfrak{h}},g^{n}\operatorname{\mathfrak{h}},g^{2n}\operatorname{\mathfrak{h}}\}$ is a facing triple. Thus if $g$ does not skewer $\operatorname{\mathfrak{h}}$ then it must be parallel to it. ∎
@@
We prove a slightly more general statement to that of Theorem 1.1:
Theorem 3.1.
Let $G$ be a finitely generated group. Suppose that $H\leq G$ is a finitely generated subgroup such that $H\operatorname{\setminus}G$ is quasi-isometric to a tree with $2\leq e<\infty$ ends. Then the following are equivalent:
  1. 1.
    $H$ is a virtual fiber subgroup.
  2. 2.
    $H$ is separable.
  3. 3.
    $H$ has bounded packing.
  4. 4.
    There are infinitely many distinct double cosets of $H$.
The following is an immediate consequence.
Corollary 3.2.
Let $H\leq G$ be a finitely generated pair of groups and suppose that $H\operatorname{\setminus}G$ is quasi-isometric to a tree with $3\leq e<\infty$ ends. Then $H$ has finitely many double cosets in $G$.
@@
The following is essentially Proposition 3.15 in [AG23], following the observation that the proof does not require $H\operatorname{\setminus}G$ to be two-ended.
@@
Let $G$ be finitely generated and let $H\leq K\leq G$ be a tower of subgroups such that $H\operatorname{\setminus}G$ is quasi-isometric to a finitely ended tree. Then either $[K:H]<\infty$ or $[G:K]<\infty$.
@@
Let $X$ be the Cayley graph of $G$ with respect to a finite generating set $S$. Since $H\operatorname{\setminus}G$ is quasi-isometric to a tree, there is a constant $D\geq 0$ such that any ball of radius $D$ separates $X$ into at least two unbounded components (this will play the role of $C_{1}$ in [AG23]). Next observe that the proof of [AG23, Lemma 3.14] uses only this property of quasi-lines so the conclusion still holds whenever $X$ a quasi-tree. The proposition therefore follows from the proof of [AG23, Proposition 3.15], subject to choosing (using the notation of that proof) $u_{1},\dots,u_{e}\in H\operatorname{\setminus}X$ such that, for each end $\xi$ of $H\operatorname{\setminus}X$, there is some $i$ such that the lift of the geodesic ray $\overline{\delta}\subseteq K\operatorname{\setminus}X$ starting at $u_{i}$ lies in $\xi$. ∎
Lemma 3.4.
Let $G$ be a group with finite generating set $S$ and $H\leq G$ be a finitely generated subgroup such that $X_{H}$ is quasi-isometric to a tree with $2\leq e<\infty$ ends. Suppose that $K\leq G$ is finitely generated. Then there exists $N\in\mathbb{N}$ such that, for any $g_{1},\dots,g_{N+1}\in G$, either
  • $\rho_{H}(g_{i}K)=\rho_{H}(g_{j}K)$ for some $i\neq j$ or
  • $\rho_{H}(g_{i}K)$ is bounded for some $i$.
@@
First note that, for any $g_{1},g_{2}\in G$, if $\rho_{H}(g_{1}K)\cap\rho_{H}(g_{2}K)\neq\emptyset$ then $\rho_{H}(g_{1}K)=\rho_{H}(g_{2}K)$.
Let $D\geq 0$ be such that any ball of radius $D$ separates $X_{H}$ into at least two unbounded connected components. Fix a finite generating set $T$ for $K$ and suppose in addition that $D>|t|_{S}$ for each $t\in T$, where |.$|.|_{S}$ denotes the word length of an element with respect to $S\cup S^{-1}$. Note that $\rho_{H}(gK)$ is $D$-coarsely connected for all $g\in G$. Let $R\geq 1$ be such that any ball of radius $D$ in $X_{H}$ contains at most $R$ vertices and let $N\coloneqq eR$. Let $\xi_{1},\dots,\xi_{e}$ denote the ends of $X_{H}$. Choose $g_{1},\dots,g_{N+1}\in G$ such that each $\rho_{H}(g_{i}K)$ is infinite and suppose that these elements are ordered such that the following holds. There exists $21=N_{0}\leq N_{1}\leq\dots\leq N_{e-1}\leq N_{e}=N+$ 2 such that, for each $i\in\{1,\dots,e\}$, if $j\in[N_{i-1},N_{i})$, then there is a subset of $\rho_{H}(g_{j}K)$ which converges towards $\xi_{i}$. It follows that, for each $i\in\{1,\dots,e\}$, there is a ball $B_{i}\subseteq X_{H}$ of radius $D$ which intersects $\rho_{H}(g_{j}K)$ for all $j\in[N_{i-1},N_{i})$ Since the number of the vertices in $\cup_{i=1}^{e}B_{i}$ is at most $N$, there exists $j_{1}\neq j_{2}$ such that $\rho_{H}(g_{j_{1}}K)\cap\rho_{H}(g_{j_{2}}K)\neq\emptyset$ and therefore $\rho_{H}(g_{j_{1}}K)=\rho_{H}(g_{j_{2}}K)$. ∎
Proposition 3.5.
Let $G$ be a group with finite generating set $S$ and $H\leq G$ be a finitely generated subgroup such that $H\operatorname{\setminus}G$ is quasi-isometric to a tree with $2\leq e<\infty$ ends. Then either the set of double cosets $\{HgH:g\in G\}$ is finite or $H$ has bounded packing in $G$.
Proof.
By [Connor-Mihalik-2013, Corollary 2.4], an element $g\in G$ is in the commensurator $\operatorname{Comm}_{G}(H)$ if and only if the cosets $H$ and $gH$ are within finite Hausdorff distance from each other. Thus $g\in\operatorname{Comm}_{G}(H)$ if and only if $\rho_{H}(gH)$ is bounded. By Lemma 3.4, there exists $N\in\mathbb{N}$ and $g_{1},\dots,g_{N}\in G$ such that, for all $g\in G-\operatorname{Comm}_{G}(H)$, there exists $i\in\{1,\dots,N\}$ such that $HgH=Hg_{i}H$. Therefore if the set $\{HgH:g\in G\}$ is infinite, then there is an infinite family of double cosets $\{Hk_{i}H\}_{i\in\mathbb{N}}$ such that each $k_{i}\in\operatorname{Comm}_{G}(H)$. Thus $H$ has infinite index in $\operatorname{Comm}_{G}(H)$. This implies that $\operatorname{Comm}_{G}(H)$ has finite index in $G$ by Proposition 3.3. By [Connor-Mihalik-2013, Lemma 6.1], $H$ has bounded packing in $\operatorname{Comm}_{G}(H)$ and by [Hruska-Wise, Proposition 2.5] this implies that $H$ has bounded packing in $G$. ∎
Proposition 3.6.
Let $G$ be a group with finite generating set $S$ and $H\leq G$ be a finitely generated subgroup such that $H\operatorname{\setminus}G$ is quasi-isometric to a tree with $2\leq e<\infty$ ends. If $H$ has bounded packing then $H$ is a virtual fiber subgroup.
Proof.
@@
Since $H$ has at least two relative ends and bounded packing, there is an action of $G$ on an unbounded finite dimensional CAT(0) cube complex $C$ with a single orbit of hyperplanes $G\cdot\operatorname{\mathfrak{h}}$ such that the $G$-stabiliser of $\operatorname{\mathfrak{h}}$, denoted $\operatorname{Stab}_{G}(\operatorname{\mathfrak{h}})$, is commensurable with $H$ (see Section 2). We assume without loss of generality that $H=\operatorname{Stab}_{G}(\operatorname{\mathfrak{h}})$. By Proposition 2.1, the action of $G$ on $C$ is essential and there is a hyperbolic element $\gamma\in G$ which skewers $\mathfrak{h}$. Label the halfspaces of $\operatorname{\mathfrak{h}}$ and replace $\gamma$ with a proper power if necessary so that $\gamma\operatorname{\mathfrak{h}}^{+}\subsetneq\operatorname{\mathfrak{h}}^{+}$. No power of $\gamma$ is contained in $H$ so $\rho_{H}(\operatorname{\langle}\gamma\operatorname{\rangle})$ is infinite. By Lemma 3.4, there exists $N\in\mathbb{N}$ and $k_{1},\dots,k_{N}\in G$ such that the $\rho_{H}$-images of $\{k_{i}\operatorname{\langle}\gamma\operatorname{\rangle}\}$ are infinite and pairwise distinct and, for all $k\in G$, if $\rho_{H}(k\operatorname{\langle}\gamma\operatorname{\rangle})$ is infinite then $\rho_{H}(k\operatorname{\langle}\gamma\operatorname{\rangle})=\rho_{H}(k_{i}\operatorname{\langle}\gamma\operatorname{\rangle})$ for some $i$.
@@ 4 Reducing a Euclidean cube complex to a line
It follows from [CFI, Lemma 2.24] that $d_{Haus}(\operatorname{\mathfrak{h}},\gamma\operatorname{\mathfrak{h}})\leq|\mathcal{H}(\operatorname{\mathfrak{h}})\Delta\mathcal{H}(\gamma\operatorname{\mathfrak{h}})|+d(\operatorname{\mathfrak{h}},\gamma\operatorname{\mathfrak{h}})$ therefore $d_{Haus}(\operatorname{\mathfrak{h}},\gamma\operatorname{\mathfrak{h}})<\infty$. ∎
@@
Let $M^{+}$ (respectively $M^{-}$) be the set of hyperplanes of $Y$ which lie in $\operatorname{\mathfrak{h}}^{+}$ (respectively $\operatorname{\mathfrak{h}}^{-}$). Then $kM^{+}=M^{+}$ for any $k\in K$. Moreover, $\gamma M^{+}\subseteq M^{+}$ and, if $\operatorname{\mathfrak{k}}\in M^{+}-\gamma M^{+}$, then $\operatorname{\mathfrak{k}}$ either separates $\gamma\operatorname{\mathfrak{h}}$ from $\operatorname{\mathfrak{h}}$ or $\operatorname{\mathfrak{k}}$ intersects $\gamma\operatorname{\mathfrak{h}}$ but not $\operatorname{\mathfrak{h}}$, so $|M^{+}-\gamma M^{+}|<\infty$ by Lemma 4.1 and thus $|M^{+}\Delta\gamma M^{+}|<\infty$. Similarly, $|M^{+}\Delta\gamma^{-1}M^{+}|=|\gamma^{-1}M^{+}-M^{+}|<\infty$. If $N$ is a set of hyperplanes such that $|M^{+}\Delta N|<\infty$ then $|M^{+}\Delta\gamma^{\pm 1}N|\leq|M^{+}\Delta\gamma^{\pm 1}M^{+}|+|\gamma^{\pm 1}(M^{+}\Delta N)|<\infty$ and, if $k\in K$, then $|M^{+}\Delta kN|\leq|M^{+}\Delta kM^{+}|+|k(M^{+}\Delta N)|<\infty$. It follows that $|M^{+}\Delta gM^{+}|<\infty$ for all $g\in\Gamma^{\prime}$, so $M^{+}$ is commensurated by the action of Γsuperscriptnormal-$\Gamma^{\prime}$. A symmetric argument shows that $M^{-}$ is also commensurated by the action of $\Gamma^{\prime}$. This implies that the map $tr_{M^{+}}:\Gamma^{\prime}\rightarrow\mathbb{Z}$ given by
@@
is well defined and, by [Cornulier, Proposition 4.H.1], it is a homomorphism.
@@
If we assume, as in the proof of Proposition 3.6, that $\Gamma$ and $\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}})$ are finitely generated and $\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}})\operatorname{\setminus}\Gamma$ is quasi-isometric to a tree with $2\leq e<\infty$ ends, then there is a much shorter proof:
For each $n\in\mathbb{N}$ let $\mathcal{H}_{n}$ denote the set of hyperplanes whose distance from $\operatorname{\mathfrak{h}}$ is exactly $n$. Then the hyperplanes of $\mathcal{H}_{n}$ contained in $\operatorname{\mathfrak{h}}^{-}$ are pairwise intersecting, as are those contained in $\operatorname{\mathfrak{h}}^{+}$. Therefore $|\mathcal{H}_{n}|\leq 2\dim(Y)$ and, since $\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}})$ acts on $\mathcal{H}_{n}$, there is a homomorphism $f_{n}:\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}})\rightarrow S(2\dim(C))$, where $S(2\dim(C))$ is the symmetric group on $2\dim(C)$ elements and $\ker f_{n}$ acts trivially on $\mathcal{H}_{n}$. The kernel $\ker f_{n}$ has index at most $(2\dim(C))!$ in $\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}})$ and, since $\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}})$ is finitely generated, there are finitely many such subgroups. Thus the intersection $K\coloneqq\cap_{n\in\mathbb{N}}\ker f_{n}$ is a finite intersection of finite index subgroups of $\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}})$ and is itself a subgroup of $\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}})$ with finite index. Let $g\in\Gamma$ be such that $g\operatorname{\mathfrak{h}}^{+}\subsetneq\operatorname{\mathfrak{h}}^{+}$. Then $K$ fixes $g^{n}\operatorname{\mathfrak{h}}$ for all $n\in\mathbb{Z}$, so $K$ has infinite index in its normaliser, which implies that $[\Gamma:N(K)]<\infty$ by Proposition 3.3. The quotient $N(K)/K$ has $e$ ends but, since it is a group, $e$ must equal $2$ so $N(K)/K$ is virtually cyclic [Hopf].

### v2

@@ 2 Cube complexes and facing triples
First note that $C$ is proper by [Hag22, Corollary 3.9]. In particular, by [CS11, Lemma 4.3], the isometry $g$ is parallel to $\operatorname{\mathfrak{h}}$ if and only if there is a power of $g$ which stabilises $\operatorname{\mathfrak{h}}$. Moreover, it follows from Lemma 2.2 that, if $g$ is peripheral to $\operatorname{\mathfrak{h}}$, then there is some $n\in\mathbb{N}$ such that $\{\operatorname{\mathfrak{h}},g^{n}\operatorname{\mathfrak{h}},g^{2n}\operatorname{\mathfrak{h}}\}$ is a facing triple. Thus if $g$ does not skewer $\operatorname{\mathfrak{h}}$ then it must be parallel to it. ∎
3 Narrow spaces
Given a connected graph $Y$, let $V(Y)$ be the vertex set of $Y$, let $d_{Y}$ (or $d$ if there is no risk of confusion) denote the path metric on $V(Y)$ where each edge has length 1 and let $B_{Y}(y,r)$ be the set of vertices with distance at most $r$ from a vertex $y\in V(Y)$. A subset $Z\subseteq V(Y)$ is said to be $\mu$-coarsely connected for some $\mu\geq 1$ if, for any $z,z^{\prime}\in Z$, there exists $\{z_{0},\dots,z_{n}\}\subseteq Z$ such that $z_{0}=z$, $z_{n}=z^{\prime}$ and $d(z_{i},z_{i+1})\leq\mu$ for each $i\in\{0,\dots,n-1\}$.
Definition 3.1.
An unbounded connected graph $Y$ is narrow if, for each $\mu\geq 1$, there exists $L(\mu)\geq 1$ such that, if $Y_{1},\dots,Y_{L(\mu)+1}\subseteq V(Y)$ are unbounded and $\mu$-coarsely connected, then there exists $i\neq j$ with $Y_{i}\cap Y_{j}\neq\emptyset$.
Proposition 3.2.
Let $Y$ be an unbounded connected graph and suppose that at least one of the following holds.
  • (a)
    $Y$ has linear growth.
  • (b)
    $Y$ is uniformly locally finite and quasi-isometric to a finitely ended tree.
Then $Y$ is narrow.
Proof.
First suppose that (a) holds and let $C\geq 1$ and $y_{0}\in V(Y)$ be such that $|B_{Y}(y_{0},n)|\leq Cn$ for all $n\in\mathbb{N}$. Fix $\mu\geq 1$ and let $L\coloneqq\mu(C+1)$. Suppose that $Y_{1},\dots,Y_{L+1}\subseteq V(Y)$ are pairwise disjoint unbounded $\mu$-coarsely connected sets of vertices, ordered such that $d(y_{0},Y_{1})\leq d(y_{0},Y_{2})\leq\dots\leq d(y_{0},Y_{L+1})$. For each $i\in\{1,\dots,L+1\}$ and $j\geq d(y_{0},Y_{i})+\mu$ we have $Y_{i}\cap(B_{Y}(y_{0},j)-B_{Y}(y_{0},j-\mu))\neq\emptyset$. Therefore if $n\geq d(y_{0},Y_{L+1})+\mu$ then $|B_{Y}(y_{0},n)-B_{Y}(y_{0},n-\mu)|\geq L+1$. Thus, if $k=d(y_{0},Y_{L+1})$, then for any $n\geq k(C+1)$
$|B_{Y}(y_{0},k+n\mu)|\geq n(L+1)\geq Ln+k(C+1)=(C+1)(k+n\mu)$
which is a contradiction.
Now suppose that (b) holds. Let $e\in\mathbb{N}$ be the number of ends of $Y$. Fix $\mu\geq 1$ and let $D\geq\mu$ be such that any ball of radius $D$ separates $Y$ into at least two unbounded connected components. Let $R\geq 1$ be such that any ball of radius $D$ in $Y$ contains at most $R$ vertices and let $L\coloneqq eR$. Let $\xi_{1},\dots,\xi_{e}$ denote the ends of $Y$. Suppose that $Y_{1},\dots,Y_{L+1}\subseteq Y$ are unbounded $\mu$-coarsely connected subspaces, ordered such that the following holds. There exists $21=L_{0}\leq L_{1}\leq\dots\leq L_{e-1}\leq L_{e}=L+$ 2 such that, for each $i\in\{1,\dots,e\}$, if $j\in[L_{i-1},L_{i})$, then there is a subset of $Y_{j}$ which converges towards $\xi_{i}$. It follows that, for each $i\in\{1,\dots,e\}$, there is a ball $B_{i}\subseteq Y$ of radius $D$ which intersects $Y_{j}$ for all $j\in[L_{i-1},L_{i})$ Since the number of the vertices in $\cup_{i=1}^{e}B_{i}$ is at most $L$, there exists $j\neq k$ such that $Y_{j}\cap Y_{k}\neq\emptyset$. ∎
Proposition 3.3.
Let $Y,Y^{\prime}$ be unbounded connected graphs. Suppose that $Y^{\prime}$ is narrow and that there exists a constant $C\geq 1$ and a metrically proper injective map $f:V(Y)\rightarrow V(Y^{\prime})$ such that $d_{Y^{\prime}}(f(x),f(y))\leq Cd_{Y}(x,y)+C$ for all $x,y\in Y$. Then $Y$ is narrow.
Proof.
For each $\mu$ let $L(\mu)$ be such that, if $Y_{1}^{\prime},\dots,Y_{L(\mu)+1}^{\prime}\subseteq V(Y^{\prime})$ are unbounded and $\mu$-coarsely connected, then there exists $i\neq j$ with $Y_{i}^{\prime}\cap Y_{j}^{\prime}\neq\emptyset$. Fix $\lambda\geq 1$ and let $Y_{1},\dots,Y_{L(C\lambda+C)+1}\subseteq V(Y)$ be unbounded and $\lambda$-coarsely connected. Then $f(Y_{1}),\dots,$$f(Y_{L(C\lambda+C)+1})\subseteq V(Y^{\prime})$ are unbounded and $(C\lambda+C)$-coarsely connected so there exists $i\neq j$ such that $f(Y_{i})\cap f(Y_{j})\neq\emptyset$. Since $f$ is injective, this implies that $Y_{i}\cap Y_{j}\neq\emptyset$. ∎
If $S,S^{\prime}$ are finite generating sets of a group $G$ and $H\leq G$ is a subgroup then the identity map $\operatorname{id}:V(H\operatorname{\setminus}\operatorname{Cay}(G,S))\rightarrow V(H\operatorname{\setminus}\operatorname{Cay}(G,S^{\prime}))$ is a bijective quasi-isometry. Therefore the above proposition implies that $H\operatorname{\setminus}\operatorname{Cay}(G,S)$ is narrow if and only if $H\operatorname{\setminus}\operatorname{Cay}(G,S^{\prime})$ is narrow. We say that $H\operatorname{\setminus}G$ is narrow if $H\operatorname{\setminus}\operatorname{Cay}(G,S)$ is narrow for some (equivalently any) finite $G$-generating set $S$.
Proposition 3.4.
Let $G$ be a finitely generated group and $H\leq G$ be such that $H\operatorname{\setminus}G$ is narrow. If $H\leq K\leq G$ is such that $[K:H]<\infty$ then $K\operatorname{\setminus}G$ is narrow.
Proof.
For any finite $G$-generating set $S$ let $X_{H,S}\coloneqq H\operatorname{\setminus}\operatorname{Cay}(G,S)$ and let $X_{K,S}\coloneqq K\operatorname{\setminus}\operatorname{Cay}(G,S)$. There is a finite degree covering map $\pi_{S}:X_{H,S}\rightarrow X_{K,S}$ and in particular any $X_{K,S}$ is unbounded. Fix a finite $G$-generating set $S$. We will show that $X_{K,S}$ is narrow. Let $\mu\geq 1$ and let $S_{\mu}\coloneqq\{g\in G:|g|_{S}\leq\mu\}$, where |.$|.|_{S}$ denotes word length with respect to $S\cup S^{-1}$. The Schreier graph $X_{H,S_{\mu}}$ is narrow so there exists $L\geq 1$ such that, for any unbounded 1-connected subsets $Z_{1},\dots,Z_{L+1}\subseteq V(X_{H,S_{\mu}})$, there exists $i\neq j$ such that $Z_{i}\cap Z_{j}\neq\emptyset$. Let $Y_{1},\dots,Y_{L+1}\subseteq V(X_{K,S})$ be unbounded and $\mu$-coarsely connected. For each $i\in\{1,\dots,L+1\}$, the subgraph of $X_{K,S_{\mu}}$ induced by $Y_{i}$ is connected so it admits a spanning tree $T_{i}$. Let $T_{i}^{\prime}\subseteq X_{H,S_{\mu}}$ be a lift of $T_{i}$ with respect to $\pi_{S_{\mu}}$. Then $V(T_{1}^{\prime}),\dots,V(T_{L+1}^{\prime})\subseteq V(X_{H,S_{\mu}})$ are unbounded and 1-connected subsets of $X_{H,S_{\mu}}$ so there exists $i\neq j$ such that there exists $x\in V(T_{i}^{\prime})\cap V(T_{j}^{\prime})$. Thus $\pi_{S_{\mu}}(x)\in V(T_{i})\cap V(T_{j})=Y_{i}\cap Y_{j}$. ∎
The following is a generalisation of Proposition 3.15 in [AG23].
Proposition 3.5.
Let $G$ be a finitely generated group and let $H\leq K\leq G$ be a tower of subgroups such that $H\operatorname{\setminus}G$ is narrow. Then either $[G:K]<\infty$ or $[K:H]<\infty$.
Proof.
Let $S$ be a finite generating set for $G$, let $X_{H}\coloneqq H\operatorname{\setminus}\operatorname{Cay}(G,S)$, $X_{K}\coloneqq K\operatorname{\setminus}\operatorname{Cay}(G,S)$ and let $\rho_{H}:\operatorname{Cay}(G,S)\rightarrow X_{H}$, $\rho_{K}:\operatorname{Cay}(G,S)\rightarrow X_{K}$ denote the quotient maps. Note that $\rho_{H}$ and $\rho_{K}$ are covers and that $\rho_{K}$ factors through a covering map $\pi:X_{H}\rightarrow X_{K}$ whose degree is $[K:H]$. Suppose that $[G:K]=\infty$, fix $\overline{x}\in X_{K}$ and let $\overline{\delta}\subseteq X_{K}$ be an infinite geodesic ray based at $\overline{x}$. For each point $x\in\pi^{-1}(\overline{x})$ there is a lift of $\overline{\delta}$ based at $x$. These lifts are pairwise disjoint and connected so $|\pi^{-1}(\overline{x})|\leq L(1)$, where $L(1)$ is as in Definition 3.1. Thus $[K:H]\leq L(1)<\infty$. ∎
@@ 3 Narrow spaces
Let $G$ be a group with finite generating set $S$ and $H\leq G$ be a finitely generated subgroup such that $X_{H}$ is narrow. Then either the set of double cosets $\{HgH:g\in G\}$ is finite or $H$ has bounded packing in $G$.
@@
By [CM13, Corollary 2.4], an element $g\in G$ is in the commensurator $\operatorname{Comm}_{G}(H)$ if and only if the cosets $H$ and $gH$ are within finite Hausdorff distance from each other. Thus $g\in\operatorname{Comm}_{G}(H)$ if and only if $\rho_{H}(gH)$ is bounded. Since $H$ is finitely generated, there exists $\mu\geq 1$ such that $\rho_{H}(gH)\subseteq X_{H}$ is $\mu$-coarsely connected for any $g\in G$. Since $X_{H}$ is narrow, there exists $N\in\mathbb{N}$ and $g_{1},\dots,g_{N}\in G$ such that, for all $g\in G-\operatorname{Comm}_{G}(H)$, there exists $i\in\{1,\dots,N\}$ such that $HgH=Hg_{i}H$. If the set $\{HgH:g\in G\}$ is infinite, then this implies that there is an infinite family of double cosets $\{Hk_{i}H\}_{i\in\mathbb{N}}$ such that $k_{i}\in\operatorname{Comm}_{G}(H)$ for each $i\in\mathbb{N}$. Thus $H$ has infinite index in $\operatorname{Comm}_{G}(H)$ so, by Proposition 3.5, $\operatorname{Comm}_{G}(H)$ has finite index in $G$. By [CM13, Lemma 6.1], $H$ has bounded packing in $\operatorname{Comm}_{G}(H)$ and by [HW09, Proposition 2.5] this implies that $H$ has bounded packing in $G$. ∎
Proposition 4.2.
Let $G$ be a group with finite generating set $S$ and $H\leq G$ be a finitely generated subgroup such that $X_{H}$ is narrow. If $H$ has bounded packing then $H$ is a virtual fiber subgroup.
@@
Since $H$ has at least two relative ends and bounded packing, there is an action of $G$ on an unbounded finite dimensional CAT(0) cube complex $C$ with a single orbit of hyperplanes $G\cdot\operatorname{\mathfrak{h}}$ such that $H$ has finite index in the $G$-stabiliser of $\operatorname{\mathfrak{h}}$, denoted $\operatorname{Stab}_{G}(\operatorname{\mathfrak{h}})$ (see Section 2). By Proposition 3.4 and since being a virtual fiber subgroup is closed under commensurability, we can assume without loss of generality that $H=\operatorname{Stab}_{G}(\operatorname{\mathfrak{h}})$. By Proposition 2.1, the action of $G$ on $C$ is essential and there is a hyperbolic element $\gamma\in G$ which skewers $\mathfrak{h}$. Label the halfspaces of $\operatorname{\mathfrak{h}}$ and replace $\gamma$ with a proper power if necessary so that $\gamma\operatorname{\mathfrak{h}}^{+}\subsetneq\operatorname{\mathfrak{h}}^{+}$. No power of $\gamma$ is contained in $H$ so $\rho_{H}(\operatorname{\langle}\gamma\operatorname{\rangle})$ is infinite. If $\mu$ is the word length of $\gamma$ with respect to $S\cup S^{-1}$ then $\rho_{H}(g\operatorname{\langle}\gamma\operatorname{\rangle})$ is $\mu$-coarsely connected for all $g\in G$. Therefore there exists $N\in\mathbb{N}$ and $k_{1},\dots,k_{N}\in G$ such that the $\rho_{H}$-images of $\{k_{i}\operatorname{\langle}\gamma\operatorname{\rangle}\}$ are infinite and pairwise distinct and, for all $k\in G$, if $\rho_{H}(k\operatorname{\langle}\gamma\operatorname{\rangle})$ is infinite then $\rho_{H}(k\operatorname{\langle}\gamma\operatorname{\rangle})=\rho_{H}(k_{i}\operatorname{\langle}\gamma\operatorname{\rangle})$ for some $i$.
@@ 5 Reducing a Euclidean cube complex to a line
It follows from [CFI16, Lemma 2.24] that $d_{Haus}(\operatorname{\mathfrak{h}},\gamma\operatorname{\mathfrak{h}})\leq|\mathcal{H}(\operatorname{\mathfrak{h}})\Delta\mathcal{H}(\gamma\operatorname{\mathfrak{h}})|+d(\operatorname{\mathfrak{h}},\gamma\operatorname{\mathfrak{h}})$ therefore $d_{Haus}(\operatorname{\mathfrak{h}},\gamma\operatorname{\mathfrak{h}})<\infty$. ∎
@@
Let $M^{+}$ (respectively $M^{-}$) be the set of hyperplanes of $Y$ which lie in $\operatorname{\mathfrak{h}}^{+}$ (respectively $\operatorname{\mathfrak{h}}^{-}$). Then $kM^{+}=M^{+}$ for any $k\in K$. Moreover, $\gamma M^{+}\subseteq M^{+}$ and, if $\operatorname{\mathfrak{k}}\in M^{+}-\gamma M^{+}$, then $\operatorname{\mathfrak{k}}$ either separates $\gamma\operatorname{\mathfrak{h}}$ from $\operatorname{\mathfrak{h}}$ or $\operatorname{\mathfrak{k}}$ intersects $\gamma\operatorname{\mathfrak{h}}$ but not $\operatorname{\mathfrak{h}}$, so $|M^{+}-\gamma M^{+}|<\infty$ by Lemma 5.1 and thus $|M^{+}\Delta\gamma M^{+}|<\infty$. Similarly, $|M^{+}\Delta\gamma^{-1}M^{+}|=|\gamma^{-1}M^{+}-M^{+}|<\infty$. If $N$ is a set of hyperplanes such that $|M^{+}\Delta N|<\infty$ then $|M^{+}\Delta\gamma^{\pm 1}N|\leq|M^{+}\Delta\gamma^{\pm 1}M^{+}|+|\gamma^{\pm 1}(M^{+}\Delta N)|<\infty$ and, if $k\in K$, then $|M^{+}\Delta kN|\leq|M^{+}\Delta kM^{+}|+|k(M^{+}\Delta N)|<\infty$. It follows that $|M^{+}\Delta gM^{+}|<\infty$ for all $g\in\Gamma^{\prime}$, so $M^{+}$ is commensurated by the action of $\Gamma^{\prime}$. A symmetric argument shows that $M^{-}$ is also commensurated by the action of $\Gamma^{\prime}$. This implies that the map $tr_{M^{+}}:\Gamma^{\prime}\rightarrow\mathbb{Z}$ given by
@@
is well defined and, by [Cor16, Proposition 4.H.1], it is a homomorphism.
@@
If we assume, as in the proof of Proposition 4.2, that $\Gamma$ and $\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}})$ are finitely generated and $\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}})\operatorname{\setminus}\Gamma$ is narrow, then there is a much shorter proof:
For each $n\in\mathbb{N}$ let $\mathcal{H}_{n}$ denote the set of hyperplanes whose distance from $\operatorname{\mathfrak{h}}$ is exactly $n$. Then the hyperplanes of $\mathcal{H}_{n}$ contained in $\operatorname{\mathfrak{h}}^{-}$ are pairwise intersecting, as are those contained in $\operatorname{\mathfrak{h}}^{+}$. Therefore $|\mathcal{H}_{n}|\leq 2\dim(Y)$ and, since $\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}})$ acts on $\mathcal{H}_{n}$, there is a homomorphism $f_{n}:\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}})\rightarrow S(2\dim(C))$, where $S(2\dim(C))$ is the symmetric group on $2\dim(C)$ elements and $\ker f_{n}$ acts trivially on $\mathcal{H}_{n}$. The kernel $\ker f_{n}$ has index at most $(2\dim(C))!$ in $\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}})$ and, since $\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}})$ is finitely generated, there are finitely many such subgroups. Thus the intersection $K\coloneqq\cap_{n\in\mathbb{N}}\ker f_{n}$ is a finite intersection of finite index normal subgroups of $\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}})$ and is itself a normal subgroup of $\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}})$ with finite index. Let $g\in\Gamma$ be such that $g\operatorname{\mathfrak{h}}^{+}\subsetneq\operatorname{\mathfrak{h}}^{+}$. Then $K$ fixes $g^{n}\operatorname{\mathfrak{h}}$ for all $n\in\mathbb{Z}$, so $K$ has infinite index in its normaliser. Since $\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}})\leq N(K)$, this implies that $[\Gamma:N(K)]<\infty$ by Proposition 3.5. The quotient group $N(K)/K$, equipped with any proper word metric, is quasi-isometric to $K\operatorname{\setminus}\Gamma$, which is a finite index cover of $\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}})\operatorname{\setminus}\Gamma$. Since narrow graphs have finitely many ends, it follows that $N(K)/K$ has $2\leq e<\infty$ ends but, since $N(K)/K$ is a group, $e$ must equal $2$ and $N(K)/K$ must be virtually cyclic [Hop44].